OpenStudy (anonymous):

Trigonometric Proofs, LAST ONES \[\frac{ \sin^4(\gamma)-\cos^4(\gamma) }{ \sin(\gamma)-\cos(\gamma)}=\sin(\gamma)\cos(\gamma)\] and \[\frac{ (1+\cos(A))(1-\cos(A))}{ \sin(A) }=\sin(A)\]

OpenStudy (anonymous):

I have no idea how to cancel out more than 1 exponent and theres no identities that match >.<

OpenStudy (misty1212):

HI!! (again again)

OpenStudy (anonymous):

xD

OpenStudy (misty1212):

more algebra i see ...

OpenStudy (misty1212):

\[\frac{b^4-a^4}{b-a}=\frac{(b^2-a^2)(b^2+a^2)}{b-a}\] is a start i.e. factor

OpenStudy (misty1212):

now we know what \(b^2+a^2\) is right ??

OpenStudy (anonymous):

Oh yes, cause that equals 1

OpenStudy (anonymous):

help misty

OpenStudy (misty1212):

ok good so we have \[\frac{b^2-a^2}{b-a}\] factor again

OpenStudy (anonymous):

so then its just left with (a^2-b^2) on top

OpenStudy (misty1212):

factor the difference of two square, cancel, have a beer

OpenStudy (anonymous):

Ah thank you. That makes so much sense.

OpenStudy (misty1212):

nice \(\gamma\) btw

OpenStudy (anonymous):

ha thank you. It was on my assignment so I assumed I should use it

OpenStudy (misty1212):

second one is real real (real) easy

OpenStudy (anonymous):

Oh yeah? am i totally missing it?

OpenStudy (misty1212):

yes you are

OpenStudy (anonymous):

misty i really need help me plzz

OpenStudy (misty1212):

\[(1-a)(1+a)=1-a^2\]

OpenStudy (misty1212):

\[\frac{1-\cos^2(\xi)}{\sin^2(\xi)}\]

OpenStudy (misty1212):

oops \[\frac{1-\cos^2(\xi)}{\sin(\xi)}\]

OpenStudy (anonymous):

that makes sense, but how do you get it to isolate sin?

OpenStudy (anonymous):

oh wait

OpenStudy (misty1212):

and guess what \(1-\cos^2(\xi)\) is equal to ?

OpenStudy (anonymous):

the indentity!

OpenStudy (misty1212):

lol

OpenStudy (anonymous):

sin^2

OpenStudy (anonymous):

then cancel ahhh yes

OpenStudy (misty1212):

cancel be done with your homework

OpenStudy (anonymous):

thank you so much, ah finally finished

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

you've been awesome!

OpenStudy (misty1212):

(blush)